Suppose that NoP(X5)6 and P(X 2):2 find P( 3< X < 4).
Answer #1

Solution:-

8) P( 3 < X < 4) = 0.0679

P(x > 5) = 0.60

x = 5

p-value for the X > 5 is = 1 - 0.60 = 0.40

z-score for the p-value = - 0.253

By applying normal distribution:-

z = rac{x-mu }{sigma }

μ-0.253 * σ 5.......................(1)

P(x < 2) = 0.20

x = 2

p-value for the X < 2 is = 0.20

z-score for the p-value = - 0.842

By applying normal distribution:-

z = rac{x-mu }{sigma }

μ -0.842 * σ 2....................(2)

By solving equation (1) and (2) we get

μ 6.29. σ 5.09

x1 = 3

x2 = 4

By applying normal distribution:-

z = rac{x-mu }{sigma }

z1 = - 0.646

z2 = - 0.449

P( - 0.647 < z < -0.449) = P(z > -0.647) - P(z > -0.449)

P( - 0.647 < z < -0.449) = 0.7412 - 0.6733

P( - 0.647 < z < -0.449) = 0.0679